OpenStudy (anonymous):

2 + (a + b) = (2 + a) + b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property @mathslover ?

mathslover (mathslover):

What is your choice unheard?

OpenStudy (anonymous):

I know it isn't Symmetric, but I'm confused on the others.

mathslover (mathslover):

Hmm, well yes its not symmetric. I have : associative property as - a + (b+c) = (a+b) + c This exactly follows : 2 + (a+b) = (2+a) + b So your answer is associative property.

OpenStudy (anonymous):

Thanks!

mathslover (mathslover):

You're welcome unheard.

OpenStudy (anonymous):

distributive property of multiplication a( x+y) = ax +ay associative property (x+y)+1= x+(y+1) commutative property of addition x+y=y+x no idea what hte symmetric property is